JET-FTIR STUDIES OF MODEL PEPTIDE SYSTEMS

\underline{COREY A. RICE} and MARTIN A. SUHM, Institut für Physikalische Chemie, Universität Göttingen, Tammannstr. 6, D-37077 Göttingen, Germany.

Small model systems, such as Formamide and N-Methylacetamide, have the model N-H⋯O=C hydrogen bond interaction, which is found in peptides and proteins. The interactions can be probed using direct absorption FTIR spectroscopy.^{a} This is a useful tool to study peptide model systems, which do not contain a UV-chromophore. Upon complexation, the aggregates show distinctive bathochromic shifts in the N-H stretching region and in the C=O stretching frequency. New developments in the coupling of FTIR spectroscopy with supersonic jet expansions for less volatile model compounds will be presented.

aRice, C. A.; Borho, N.; Suhm, M. A. Z. Phys. Chem. 219 (2005) 379–388.